OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-5, -1) and is perpendicular to y = five halves x + 2. y = two fifths x + 1 y = -five halves x - 27 halves y = -two fifths x - 3 y = five halves x + 23 halves PLEASE HELP

OpenStudy (jdoe0001):

what is the slope of \(\bf y = \cfrac{5}{2}x + 2 \quad ? \)

OpenStudy (anonymous):

5/2

OpenStudy (jdoe0001):

yes, 5/2 so a line that's perpendicular to that equation, will have a NEGATIVE RECIPROCAL slope what does that mean? well RECIPROCAL of 5/2 2/5 NEGATIVE of that -2/5 so, now we know that, the line has a slope of -2/5 and passes through the point (-5, -1) so, we'll plug those values in your "point-slope form" of \(\bf y-y_1=m(x-x_1)\) so \(\large \begin{array}{llll} &x_1&y_1&\qquad slope = -\cfrac{2}{5}\\ (&-5, & -1)\\\quad \\\quad \\ &&&\color{blue}{y-y_1=m(x-x_1)} \end{array} \)

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."